OpenStudy (anonymous):

A triangular prism has a surface area of 350 square feet, a length of 15 feet, a height of 10 feet, and a side of 5 feet. Find the width of the triangular prism Please help!

OpenStudy (anonymous):

Let's see... S. A = 350 l = 15 ft h = 10 s = 5 ft w =?  Let's flatten the prism:  We have two triangles: The area for a triangle is (w x h) / 2 Since there are two triangles it should be 2[ (w x h) / 2 ] = w x h We have three rectangles: The area for a rectangle is l x w Since there are 3 triangles, it should be 3 (l x w) So the surface area is: 3 (l x w) + (w x h) S.A = 3 (l x w) + (w x h) Solve the equation in terms of "w" S.A= 3 l + 3w + wh S.A = 3l + w( 3+h) S.A - 3l = w (3+h) w = (S.A - 3l ) / (3+h) Substitute your given values in the formula and you're done :)